OpenStudy (anonymous):

Why is the Half-Angle formula derived from the Double-Angle formula? Isn't that like ending back to where you started? Half a double? Obviously not or this wouldn't work! So can someone explain to a dummy like me? :)

OpenStudy (anonymous):

gimmick is to take \(\cos(2x)\) formula and replace \(2x\) by \(x\) and \(x\) by \(\frac{x}{2}\) then you solve for \(\cos(\frac{x}{2})\) in the resulting equation

OpenStudy (anonymous):

yes, it is a gimmick, no, you do not get back what you started because you are solving for \(\cos(\frac{x}{2})\)

OpenStudy (anonymous):

i know how to get the formula, I just don't get the why? haha :)